Progress on the 24 inch
 Overhaul - RA Dec Gears, Primary, Focus, Shutter, Filter Wheel, August 2010
 Exercise 6 for K-12 students
Asteroids, Comet, and NEO's, October 2007
 Exercise 5 for K-12 students
The Moon and Orbital Dynamics, April 2007
 Exercise 4 for K-12 students
Variable Stars, January 2007
 Steel Baffle Replacement, November 2006
 Zoom-Photo Printer Software, October, 2006
 Exercise 2 and 3, for K-12 students
Scientific Method and The Blue Sky on-line, September 2006
 Telescope Overhaul and New Filters, August 2006
 IDEAS Grant Meeting in Silver City, July 2006
 IDEAS Grant Awarded to UALR for NFO Webscope, June 2006
 Correlating Lightning Detectors, May 2006
 Replacing the Dec Worm Bearings, April 2006
 Photometric Corrections, March 2006
 Deep Sky Field NFO3, November 2005
 Just a regular Night at Work, September 2005
 Local lightning activity added to weather data, August 2005
 Exercise 1, for K-12 students, Tri-color Image on-line, August 2005
 Backup Sensors, May 2005
 World Coordinate Data(WCS) added to FITS headers, April 2005
 Gravity Anti-lash on RA, April 2005
 Conference at UALR Feb. 2005
 Image Requests and Dispersals Completely Automated, 01/24/05
 Scientific Images on Asteroids, Supernova Searches and Eclipsing Binaries Underway 12/21/04
data1 data2 data3
 Drive Gears Upgraded, 11/04
 True Color Combined Images M27 M16 M17 M57 NGC6940 NGC7009 09/28/04
 False Color Images M27 M16 M17 M57 M33 and NGC253 Trifid 09/28/04
 Image 'pipeline' to in Arkansas enabled, 09/20/04
 RA drive belt upsized, 09/15/04
 The worm drive tracking is corrected, Dumbell Nebula 08/25/04
 Images Registered 'NSEW' A B 08/19/04
 Images Improving A B C D 08/18/04
 First light for the camera, M5, 08/08/04
 Mounting and Balancing the Camera on the telescope, 08/03/04
 Meeting with teachers from Silver Schools, 8/03/04
 More work on the New CCD Camera 06/04
 The New CCD Camera, Final Assembly 04/26/04
 The New CCD Camera 03/15/04
 Main Weather Station on-line, February 7, 2004 from local network
 Cloud and Acute Moisture Sensors, February 2004
 Buffer for Interface to the Filter Wheel and Shutter, January 2004
 Work Progressing on the Main Instrument, January 2004
 The Weather Station is Installed, January 2004
 Dr. Lacy works on Pointing and Dr. Grauer, Shaw and McGuire visit from Arkansas November, 2003
 The computers move to the Dome 10/10/03
 First Light Celebration, Target Mars 09/27/03
 Now Just Software Remains 09/26/03
 Mirror back from Recoating and Installed 09/21/03
 Ready to install the Optics 09/14/03
 Rough Polar Alignment (about <10 arc minutes) 9/12/03
 Re-greasing Dec and RA and Mounting the Base on the Pedestal 09/10/03
 Re-greasing, repainting and Completion of the Digital Focus 09/01/03
 Removal of the Dec Bearing for Regreasing 08/22/03
 Mirror Removed for Recoating and Focus Electronics 7/27/03
 Lightning Protection, Grounding, Computer Installed, Dome On-Line 5/24/03
 Roll-Off Roof Fully Functional 4/20/03
 Dome Motor Control and Computer Interface 4/4/03
 Exterior Finished and Painted 03/12/03
 Crane Day 02/28/03
 Dome Painted and Ready for the crane 2/15/03
 Choose the new logo    EMAIL Bill at
 Dome assembly Day 2 02/02/03
 Dome assembly Day 1 01/25/03
 Instrument Box, Sandblasting and Painting 01/07/03
 The Roll-Off Carier is Installed 11/23/02
 More Steel Stucture and the trip for the Colloquium 10/11/02
 The Steel Stucture is completed 9/05/02
 The Cement is Poured 8/30/02
 Power is connected and the Stucture Grows 8/20/02
 Corner Post, Power and Steel Delivered 8/10/02
 Power Pole Set and Footers Dug 7/30/02
 Dome Construction Drawings 7/02
 Position sensors and anti-lash for DEC, RA, and Worm 6/02
 Drive Motors and Belts 5/02
 Assembly of the scope at the shop 4/02
 Rolling Mount ready for the telescope 3/02
 Progress on the Tracking Program 2/02
 Trip to Arkansas to pick up the scope 8/01